KyledaGreat:

5x + 8y = 4y − 5 Find the equation of the line which passes through the point (11,12) and is perpendicular to the given line. Express your answer in slope-intercept form. Simplify your answer.

CortneyTubbs:

do you know what the slope intercept form is?

KyledaGreat:

yes

CortneyTubbs:

ok so do you know how to get that equation to slope intercept form?

KyledaGreat:

no

CortneyTubbs:

ok so u need to get the y by itself

KyledaGreat:

you mean like this y =

CortneyTubbs:

yes

CortneyTubbs:

do you know how you would do that?

KyledaGreat:

no indeed

CortneyTubbs:

ok so you do the following:\[5x-5x+8y=4y-5x-5\]\[8y-4y=4y-4y-5x-5\]\[4y=-5x-5\]

CortneyTubbs:

I got you to there you need to divide both sides by 4 but first do you get how i got 4y=-5x-5?

KyledaGreat:

not exactly per se

KyledaGreat:

and how do i divide both sides by 4, i'm sorry i wasn't all that good with math

CortneyTubbs:

its ok. the first thing i did was subtract 5x from both sides. Then i subtracted 4y from both sides.

CortneyTubbs:

do you get that?

KyledaGreat:

yes sort of , i'm just trying to figure out how to divide from both sides in simpler times for me to get it

KyledaGreat:

yes sort of , i'm just trying to figure out how to divide from both sides in simpler times for me to get it

KyledaGreat:

terms i meant

CortneyTubbs:

so one side has 4y and we need to get the y by itself right?

KyledaGreat:

yes

CortneyTubbs:

ok so you divide 4y by 4 to get y all by itself

CortneyTubbs:

but what you do to one side of the equal sign you have to do to the other side

CortneyTubbs:

so you take -5x-5 and divide it by 4

KyledaGreat:

okay , is it -5x - 5/4 ?

CortneyTubbs:

so the equation now would be y=-1.75x-1.75

KyledaGreat:

-1.75

CortneyTubbs:

ok but it wants the perpendicular equation that lines up with (11,12)

KyledaGreat:

i see , how could i get it up to there ?

KyledaGreat:

i think i might be right or wrong , is it y = 4/5x + 16/5 ?

CortneyTubbs:

i am trying to figure that out rn. its been a while since i have done these things

KyledaGreat:

i understand , i've been out of school since may as i graduated and all this bring me back so i get what you mean lol

CortneyTubbs:

I am a super senior in high school

CortneyTubbs:

how did you get that equation?

KyledaGreat:

basically, i looked back and i know that it would have to be y - y1 = m(x - x1) to find the line perpendicular

CortneyTubbs:

that is the correct answer but I have no clue how you got it.

KyledaGreat:

thank you , i seen from an example . i'll show you

CortneyTubbs:

ok

KyledaGreat:

1 attachment